# Natural Language Processing RELIES on Linguistics

### v3

@@ Abstract
Large Language Models (LLMs) have become capable of generating highly fluent text in certain languages, without modules specially designed to capture grammar or semantic coherence. What does this mean for the future of linguistic expertise in NLP? We highlight several aspects in which NLP (still) relies on linguistics, or where linguistic thinking can illuminate new directions. We argue our case around the acronym RELIES that encapsulates six major facets where linguistics contributes to NLP: Resources, Evaluation, Low-resource settings, Interpretability, Explanation, and the Study of language. This list is not exhaustive, nor is linguistics the main point of reference for every effort under these themes; but at a macro level, these facets highlight the enduring importance of studying machine systems vis-à-vis systems of human language.
@@ 1 Introduction
It is 2024. ChatGPT has been an international sensation for over a year, a focal point in the ongoing public hype, enthusiasm, and concern about “generative AI.” NLP conferences are awash in papers about prompting Large Language Models that have gobbled zillions of words of text and can produce eerily fluent language (in English, anyway). These models have been refined with corrective feedback on conversational outputs, but (so far as we know) no direct inductive biases about natural language grammaticality or compositional meaning. [^1] For example, we did not find any such linguistic biases in the reported Llama 3.1 training procedure, though parts of the training targeted code (Llama Team, 2024). Like BERT and company an NLP research epoch before—or word embeddings an epoch before that—they have elicited both fascination and hand-wringing in scientific circles.
@@
1.1 What do we mean by “linguistics”, “NLP”, and “CL”?
Before developing our argument, we need to define our terms (while acknowledging that our definitions will not be perfectly crisp, as disciplinary boundaries are inherently fuzzy).
Linguistics is the study of systematicity and variation in communication between humans, as transmitted via speech, sign, and writing. Linguists may specialize in how knowledge of forms is organized and deployed to symbolize meanings; how linguistic behaviors emerge, drawing on human cognitive and social capacities; and how language varies over time (ontogeny, phylogeny), geography, identity group, etc.
Adjacent to the core field of linguistics are a range of fields of study and practice that center on language, including language education, translation, documentation, rhetoric, communications, and philosophy of language. Below, the term “linguistics” covers these (non-AI) “language fields”, broadly defined.
Natural Language Processing (NLP) is the field concerned with developing technology for sophisticated computational processing of text, and especially, computational understanding or generation of individual sentences, documents, or conversations (as opposed to drawing inferences about entire collections). [^3] Text mining, information retrieval, and speech processing are studied in separate research communities, though there is some overlap, as evidenced by dedicated tracks in NLP conferences such as ACL. These area, in combination with prototypical NLP applications such as machine translation and summarization, constitute language technologies. The shifting technological landscape may be opening up opportunities for greater unification of these fields (e.g., Chrupała, 2023). Contemporary NLP research is organized around a paradigm of empirical evaluation of systems on tasks, whether connected with general user applications (like question answering, translation, summarization) or more granular and focused on aspects of the language system (like parsing and coreference resolution). In recent years, the dominant concern of NLP has been the design, interpretation, and application of pretrained transformer language models such as BERT and GPT. Pretrained models of considerable scale are often called Large Language Models (LLMs).
@@ 1.1 What do we mean by “linguistics”, “NLP”, and “CL”?
1.2 Where would NLP be without linguistics?
Linguistics is no longer front and center in the way we build NLP systems for practical tasks (§2). Gone are the days when machine translation engines consisted of rules painstakingly crafted by linguists. But engineering the behavior of a system is only one component of the NLP research or development/deployment lifecycle, and there are ways expertise in linguistics continues to be essential in NLP—that is to say, there are ways NLP relies on linguistics.
Imagine what the field would be like without linguistics, assuming for the sake of argument that current modeling approaches (word embeddings, transformers, LLMs) had somehow been developed with only a cursory understanding of how language works. Many of the systems might look similar, but the field would look very different. The field would have only the shallowest attention to linguistic analysis (i.e. tasks that focus on representing regularities in linguistic systems beyond the information communicated in a particular application setting). Even for objectives not primarily tied to linguistic analysis, our toolbox would be all the poorer. We explore six facets of NLP research, organized under the mnemonic “RELIES.” Indeed, without linguistics
@@ 1.2 Where would NLP be without linguistics?
  • Study of language: Classic cL tasks such as parsing, coreference resolution, and textual entailment/natural language inference would not exist within our community. An NLP devoted purely to commercial technology would also be indifferent to applications connected with scholarly or community-driven linguistic work, such as documenting endangered languages.
In what follows, we give a sampling of the ways linguistics continues to play a role in NLP, organized under the RELIES categories listed above: resources (§3), evaluation (§4), low-resource settings (§5), interpretability/explainability (§6), and study of language (§7). It is worth emphasizing that these categories and the literature review are meant to be illustrative, not exhaustive; that we are speaking at a macro level; and that we are not claiming that linguistics is the only or necessarily the single most important source of expertise for working with language data and systems. [^4] That is, the RELIES facets highlight ways that linguistics can be useful in individual projects, and has a role to play in the field at large, alongside other kinds of expertise. We are not claiming that advanced training in linguistics is always a prerequisite for engaging with the RELIES facets (it depends on the nature of the study). We conclude by discussing pertinent challenges for the NLP community which involve linguistic expertise (§8).
@@ 2 Background
Today it is not a given that theories and representations from linguistics will form a direct foundation of NLP technologies. Given the current state of the field, it is possible to do research in NLP without traditional training in linguistics. As was the case in the age of the “statistical revolution” of NLP (Johnson, 2009; Uszkoreit, 2009; Church, 2011), the neural era has brought on skepticism towards the role of linguistics in NLP engineering tasks, offering ways to do language processing ‘from scratch’ Collobert et al. (2011). The field also saw the introduction of pre-trained word embeddings Mikolov et al. (2013) and then pre-trained contextualized language models Howard and Ruder (2018), with the well-known inflection points of ELMo and BERT Peters et al. (2018); Devlin et al. (2019). With such milestones as ChatGPT Open AI (2022) in 2022/23, the past couple of years have seen LLMs with billions of parameters that can be accessed by a wide range of users through a natural language interface (‘prompting’).
Do these developments prove the ‘bitter lesson’ Sutton (2019) that AI for language is best achieved by funneling data on a massive scale into rather general machine learning architectures? What role does linguistics play nowadays in NLP? Triggered by these and similar questions, we analyze the role of linguistics in the current LLM regime of NLP.
We are not the only researchers that have studied these apparently tectonic shifts. Relatedly, Ignat et al. (2024) collect interesting topics for PhD students, particularly evaluation, and Saphra et al. (2023) hypothesize a cyclical historic model that then would suggest that familiar problems will resurface.
The relationship between Linguistics and NLP (as well as the relationship between Linguistics and AI) has been a topic of discussion for some time (e.g., Lakoff, 1978; Raskin, 1985; Nirenburg, 1986). Within the ACL community, it has been explicitly on the agenda of fora such as the EACL 2009 Workshop on the Interaction between Linguistics and Computational Linguistics: Virtuous, Vicious or Vacuous? (Baldwin and Kordoni, 2009), the ACL 2010 Workshop on NLP and Linguistics: Finding the Common Ground (Xia et al., 2010), and more recently the NAACL 2022 panel entitled “The Place of Linguistics and Symbolic Structures” (Carpuat et al., 2022, p. xxxiv).
In this paper, we want to look more generally at the overlap between the concerns of linguistics and the concerns of contemporary NLP as well as their relationship. In the sections that follow, we tackle each of the six facets of NLP research which make up the mnemonic “RELIES,” for which we argue that linguistics has enduring relevance.
@@ 3 Resources
The field of NLP is committed to an empirical methodology wherein machine learning models are trained and evaluated on language data. This paradigm requires linguistic expertise on several fronts. Resources are supported by various degrees of linguistic knowledge—ranging from proficiency in a language to formal training in linguistics.
Creation of resources for general NLP tasks
@@
By “general NLP tasks” we mean tasks that closely relate to applications in widespread demand, such as machine translation (MT), entity linking, and sentiment classification. Studying these tasks in an empirical way requires corpus resources. Even if we can do without some of these resources in the training of NLP systems, they remain relevant for testing and studying systems, ensuring sound behavior.
Linguistic expertise contributes to the development of resources at the first step of building a dataset: selecting data. The selection and curation of data is often informed by language expertise, even if the data is raw text.
Linguistic expertise is important in moving away from a language ideology which sets some fluent text as more “high quality” than others and ripe for inclusion in training data. For example, awareness of various sociolinguistic factors can inhibit bias in model training and use: Gururangan et al. (2022) showed that US school newspaper data is more likely to be marked as “high quality” via GPT-3 if it comes from a school in a wealthier zip code.
Issues that arise in selection/curation of raw text include language diversity, variation amongst dialects/speakers, genre variation, code-switching, social biases, and undesirable content (e.g. hate speech). This is particularly important because of the wide variability of human language and the subjectivity of core NLP tasks; for example, Sasidharan Nair et al. (2024) explores the reliability and reproducibility of sentiment annotations of code-mixed data, and Abercrombie et al. (2023) investigates the impact of source language on hate speech labeling.
Data selection may involve shallow linguistic heuristics such as the analysis of word/sentence counts, or linguistically advanced techniques that increase and validate the diversity and language phenomena coverage of the data Dryer (1989); Rijkhoff et al. (1993). For example, Ravichander et al. (2022) select data with sufficient prepositions and complementizers, and Ponti et al. (2020) inform their data selection with a language typology index Littell et al. (2017). Statistical induction of dataset properties is possible through assessing its broader linguistic features, as shown by (Monteiro et al., 2024) that characterize datasets with psycholinguistic features such as readability index (Anderson, 1983) or content word frequency (Baayen et al., 1995).
@@
In the annotation process, advanced linguistic knowledge of annotators can also be important. For example, people skilled at language analysis are needed to ensure meaningful evaluation of machine translation systems. As such, paraphrasing references with linguists increases their quality and diversity (Freitag et al., 2020), and professional translators (Freitag et al., 2021a) offer annotations that provide correct system rankings (in both studies, crowd-workers were not able to make the cut). For consistent annotation of sentiment, Taboada (2016) outline that it is necessary to understand and conceptualize linguistic phenomena, such as discourse patterns. Since not all tasks may be suitable for crowdsourcing without extensive experimentation, linguistically trained annotators can also help reduce annotation cost (Gillick and Liu, 2010).
Further, the design and execution of crowdsourcing and annotation protocols depend on (advanced) linguistic knowledge of designers. This is due to the challenging nature of establishing communication that will elicit accurate judgments from lay people facing diverse natural language data Ambati et al. (2012); an agreed-upon linguistic metalanguage and vocabulary is required to invoke a consensual view on the setting and the linguistic object of interest. This holds true even for tasks such as entity linking, which though less subjective, calls for a precise definition of named entity that resolves “lexical, syntactic, semantic, and pragmatic” ambiguities Berry (2007); Bouarroudj et al. (2022). [^5] Even less subjective tasks like named entity annotation elicit annotation disagreement (Peng et al., 2024) and dataset creators and analyzers must take care to consider the range of backgrounds and perspectives which contribute to disagreement (Wein et al., 2023; Prabhakaran et al., 2024). As such, the linguistic knowledge that both or either the designer and annotator have is (either explicitly or implicitly) relevant to the creation of the dataset, even if it’s not a “linguistic” task. [^6] For more on metalanguage, see also Interpretability and Explanation (§6).
Dataset development
Annotation is a complex endeavor that with insufficient care can easily produce misleading data, and thus misleading conclusions (Tedeschi et al., 2023). This is especially true with crowdsourcing, given the difficulty of controlling the annotation quality. As such, irreproducible and low-quality crowdsourced datasets often emerge (Karpinska et al., 2021).
Therefore, each aspect of the process of language resource creation requires careful consideration (Rehm, 2016; Ide and Pustejovsky, 2017; Klie et al., 2024), including the development and documentation of data. One step towards consistent data use is standardized documentation of the characteristics and intended uses of datasets (Gebru et al., 2021; McMillan-Major et al., 2024).

### v4

@@ Abstract
Large Language Models (LLMs) have become capable of generating highly fluent text in certain languages, without modules specially designed to capture grammar or semantic coherence. What does this mean for the future of linguistic expertise in NLP? We highlight several aspects in which NLP (still) relies on linguistics, or where linguistic thinking can illuminate new directions. We argue our case around the acronym RELIES, which encapsulates six major facets where linguistics contributes to NLP: Resources, Evaluation, Low-resource settings, Interpretability, Explanation, and the Study of language. This list is not exhaustive, nor is linguistics the main point of reference for every effort under these themes; but at a macro level, these facets highlight the enduring importance of studying machine systems vis-à-vis systems of human language.
@@ 1 Introduction
It is 2025. ChatGPT has been an international sensation for over two years, a focal point in the ongoing public hype, enthusiasm, and concern about “generative AI.” NLP conferences are awash in papers about prompting Large Language Models that have gobbled zillions of words of text and can produce eerily fluent language (in English, anyway). These models have been refined with corrective feedback on conversational outputs, but (so far as we know) no direct inductive biases about grammaticality or compositional meaning. [^1] For example, we did not find any such linguistic biases in the reported Llama 3.1 training procedure, though parts of the training targeted code (Llama Team, 2024). Like BERT and company an NLP research epoch before—or word embeddings an epoch before that—they have elicited both fascination and hand-wringing in scientific circles.
@@
What do we mean by “linguistics”, “NLP”, and “CL”? Before developing our argument, we need to define our terms (while acknowledging that our definitions will not be perfectly crisp, as disciplinary boundaries are inherently fuzzy).
Linguistics is the study of systematicity and variation in human communication, as transmitted via speech, sign, and writing. Linguists may specialize in how knowledge of forms is organized and deployed to symbolize meanings; how linguistic behaviors emerge, drawing on human cognitive and social capacities; and how language varies over time (ontogeny, phylogeny), geography, identity group, etc. Linguistics-adjacent fields of language study and practice include language education, translation, rhetoric, communications, and philosophy of language. Below, the term “linguistics” covers these (non-AI) “language fields”, broadly defined.
Natural Language Processing (NLP) is the field concerned with developing technology for sophisticated computational processing of text, and especially, computational understanding or generation of individual sentences, documents, or conversations (as opposed to drawing inferences about entire collections). [^3] Text mining, information retrieval, and speech processing are studied in separate research communities, with some overlap. These areas, in combination with prototypical NLP applications such as machine translation and summarization, constitute language technologies. The shifting technological landscape may be opening up opportunities for greater unification of these fields (e.g., Chrupała, 2023). Contemporary NLP research prioritizes empirical evaluation of systems on tasks, whether connected with general user applications (like QA, translation, summarization) or more granular and focused on aspects of the language system (like parsing and coreference resolution). In recent years, the dominant concern of NLP has been the design, interpretation, and application of pretrained transformer language models such as BERT and GPT. Pretrained models of considerable scale are often called Large Language Models (LLMs).
@@
Where would NLP be without linguistics? Linguistics is no longer front and center in the way we build NLP systems for practical tasks. Gone are the days when machine translation engines consisted of rules painstakingly crafted by linguists. But there are ways expertise in linguistics continues to be essential in NLP—that is to say, there are ways NLP relies on linguistics.
Imagine what the field would be like without linguistics, assuming for the sake of argument that current modeling approaches (word embeddings, transformers, LLMs) had somehow been developed with only a cursory understanding of how language works. Many of the systems might look similar, but the field would look very different. The field would have only the shallowest attention to linguistic analysis (i.e. tasks that focus on representing regularities in linguistic systems beyond the information communicated in a particular application setting). Even for objectives not primarily tied to linguistic analysis, our toolbox would be all the poorer:
@@
  • Study of language: Classic cL tasks such as parsing, coreference resolution, and textual entailment would not exist within our community. An NLP devoted purely to commercial technology would also be indifferent to goals of scholarly or community-driven linguistic work, such as documenting endangered languages.
In what follows, we give a sampling of the ways linguistics continues to play a role in NLP, organized under the mnemonic “RELIES”: resources (section 3), evaluation (section 4), low-resource settings (section 5), interpretability/explanation (section 6), and study of language (section 7). It is worth emphasizing that these categories and the literature review are meant to be illustrative, not exhaustive; that we are speaking at a macro level; and that we are not claiming that linguistics is the only or necessarily the single most important source of expertise for working with language data and systems. We conclude by discussing pertinent challenges for the NLP community which involve linguistic expertise (section 8).
@@ 2 Background
Today it is not a given that theories and representations from linguistics will form a direct foundation of NLP technologies; it is possible to do research in NLP without traditional training in linguistics. As was the case in the age of the “statistical revolution” of NLP (Johnson, 2009; Uszkoreit, 2009; Church, 2011), deep learning was heralded as a ‘tsunami’ (Manning, 2015) and a way to model language ‘from scratch’, rather than via features reflecting linguistic expertise Collobert et al. (2011). Milestones in the neural era included the introduction of word embeddings Mikolov et al. (2013) and then pretrained language models Howard and Ruder (2018), with the well-known inflection points of ELMo and BERT Peters et al. (2018); Devlin et al. (2019) in 2018/19 and ChatGPT Open AI (2022) in 2022/23. The recent past has seen LLMs with billions of parameters that can be accessed by a wide range of users through a natural language interface (‘prompting’). Do these developments prove the ‘bitter lesson’ Sutton (2019) that language technologies are most effectively developed solely by funneling data on a massive scale into rather general machine learning architectures? What role does linguistics play nowadays in NLP?
We are not the only researchers to reflect on these apparently tectonic shifts. Ignat et al. (2024) collect interesting topics for PhD students, particularly evaluation, and Saphra et al. (2024) hypothesize a cyclical historic model that then would suggest that familiar problems will resurface. We also see that the relationship between Linguistics and NLP (as well as the relationship between Linguistics and AI) has been a topic of discussion for some time (e.g., Lakoff, 1978; Raskin, 1985; Nirenburg, 1986; Baldwin and Kordoni, 2009; Xia, Lewis, and Levin, 2010). In 2011, an entire collection of papers was devoted to the relationship between linguistics and CL (Baldwin and Kordoni, 2011). There it was noted that the field of CL encompasses both scientific and engineering perspectives (Kay, 2011; Johnson, 2011); and that even with machine learning algorithms, long-tail (rare) phenomena celebrated by linguists are likely to challenge systems (Kay, 2011; Levin, 2011; Steedman, 2011), and claims that NLP solutions will work well for any language need to be carefully evaluated in light of linguistic diversity (Bender, 2011) and low-resource conditions in small languages (Bird, 2011).
In what follows, we offer a contemporary take on how linguistics has enduring, practical relevance for NLP research, under the mnemonic “RELIES”.
@@ 3 Resources
The empirical paradigm that dominates NLP today requires language datasets for training and evaluation. Such resources are supported by various degrees of linguistic knowledge—ranging from proficiency in a language to formal training in linguistics.
Resources for general NLP tasks
@@
By “general NLP tasks” we mean tasks that closely relate to applications in widespread demand, such as machine translation (MT), summarization, and sentiment classification. Empirical study of these task requires corpus resources. Even if we could now do without some of these resources in the training of NLP systems, they remain highly relevant for testing and studying systems.
The selection and curation of data can be informed by language expertise, even if the data is raw text. We wish for reflection of language diversity, variation among dialects/speakers, genre variation, code-switching, and (if not target of a specific RQ, probably) reduced social biases or problematic content (e.g., hate speech). Sensitivity to linguistic factors is particularly important because of the wide variability of human language and the subjectivity inherent in certain labeling problems, like sentiment and toxic language (Sasidharan Nair, Dinkar, and Abercrombie, 2024; Abercrombie, Hovy, and Prabhakaran, 2023). Specific data selection techniques may involve shallow linguistic heuristics such as frequency counts, or linguistically advanced techniques that increase and validate the diversity and language phenomena coverage of the data Dryer (1989); Rijkhoff et al. (1993). For example, Ravichander, Gardner, and Marasovic (2022) select data with sufficient prepositions and complementizers, and Ponti et al. (2020) inform their data selection with a language typology index Littell et al. (2017). The BabyLM challenge (Warstadt et al., 2023; Hu et al., 2024) selects training and testing data according to cognitive linguistic plausibility, addressing fundamental questions in the learning process of LLMs.
@@
In the annotation process, advanced linguistic knowledge of annotators has been of value, and people skilled at language analysis can ensure meaningful evaluation of NLP systems. For example, Freitag, Grangier, and Caswell (2020) observe that when linguists paraphrase MT references, this increases their quality and diversity; And only professional translators (instead of crowd-workers) lead to correct MT system rankings (Freitag et al., 2021a). For consistent sentiment annotation, Taboada (2016) highlights the importance of understanding linguistic discourse phenomena. Since not all tasks may be suitable for crowdsourcing without extensive experimentation, linguistically trained annotators can help reduce annotation cost (Gillick and Liu, 2010).
Finally, designing and coordinating annotation protocols, especially via crowdsourcing, requires sensitivity to linguistic issues in order to craft guidelines and processes that produce a reasonable level of coherence across the annotations, and to recognize where disagreements can be explained by different perspectives (Ambati, Vogel, and Carbonell, 2012; Plank, 2022; Wein et al., 2023; Prabhakaran et al., 2024).
Resources with linguistic annotations
These encompass datasets of various sorts of grammatical and semantic structures. Linguistic knowledge and awareness is, of course, needed for designing analytical frameworks, and also for conducting annotation and verifying model output.
Two examples of linguistically detailed, resource-oriented frameworks prominent in current NLP research are Universal Dependencies, with morphosyntactic annotations for over 150 languages (UD; Nivre et al., 2016, 2020), and Abstract Meaning Representation, which is a framework for describing semantic graphs that has datasets in several languages (AMR; Banarescu et al., 2013; Wein and Schneider, 2024). UD and AMR are applied linguistic theories (in the sense of Raskin, 1985, p. 275). Research in and with linguistic annotation frameworks continues apace, with numerous recent workshops in the ACL community focused on grammar and semantics (Bonial and Tayyar Madabushi, 2023; Dakota et al., 2023; Bonn and Xue, 2023; Grobol and Tyers, 2023; Rambow and Lareau, 2023). The AMR framework alone is featured in 450 papers, [^4] https://nert-nlp.github.io/AMR-Bibliography/ and numerous applications (Wein and Opitz, 2024; Sadeddine, Opitz, and Suchanek, 2024).
Generally, linguistically-annotated corpora and NLP models contribute to studying questions about language (section 7). However, for some major NLP applications (like MT and QA) in high-resource settings, the future of explicit linguistic representations remains to be seen, with mixed results from neuro-symbolic models (Hamilton et al., 2022; Shwartz, 2023). On the other hand, we consider the utility of such formal representations for evaluating and interpreting models (e.g., Xu et al., 2021; Opitz and Frank, 2022; Fodor, De Deyne, and Suzuki, 2024) to be an exciting path (Evaluation, section 4; Interpretability, section 6).
4 Evaluation
We illustrate where linguistics is useful to successfully evaluate an NLP system, that is, to describe the degree to which it conforms to our expectations. For tasks that take NL input or generate NL output, it is important to measure correctness and robustness at the language-system level, for which appropriate evaluations (quantitative and qualitative) are informed by linguistic expertise. [^5] A similar sentiment was expressed by Levin (2011) in the statistical NLP era: “language technologists [ought to] understand the object of study, human language…in order to understand where current methods are falling short, we as a field need to understand the data” (p. 19).
Gold standard evaluation
Linguistic resources, particularly so-called gold standard annotations, play an important role in the evaluation of NLP technologies. They also contribute to the benchmarking of LLMs and “AI systems” (Tenney et al., 2019a; Pimentel et al., 2020a): Many of the tasks packaged in the Beyond the Imitation Game Benchmark (BIG-bench) (Srivastava et al., 2023) are linguistic tasks, such as bridging resolution and detecting common morphemes. Another kind of benchmark tests linguistic inference from just a few examples in puzzles (Şahin et al., 2020; Chi et al., 2024), a format conducive to low-resource languages without established annotation conventions (Bean et al., 2024).
Linguistic resources are also crucial for system diagnostics (including error analysis), which is a part of evaluation, but does not necessarily have the aim to produce a “global” ranking of AI systems. Instead, diagnostic methods help find more targeted improvement perspectives, with possible relevance even to broader society, since linguistic phenomena are often contingent with issues of societal interest, like gender biases in coreference (Rudinger et al., 2018) or QA (Parrish et al., 2022). [^6] Raji et al. (2021) generally argue for more targeted evaluation methods to track “AI progress.”
Overall, linguistic diagnostics can come in many forms; here we give a small sample. The ‘CheckList’ (Ribeiro et al., 2020) tests NLP systems in various aspects, using linguistic features (negation, part of speech, etc.) to build templates. Song et al. (2022) assess LLMs on Chinese linguistic phenomena, and Parcalabescu et al. (2022) employ linguistic aspects to study (computer) vision and language systems. Automated diagnostics can be achieved, e.g., through targeted measurements on parsed semantic structures (Lo and Wu, 2011; Opitz and Frank, 2021; Fan, Aumiller, and Gertz, 2023, calling for work on the intrinsic NLP task of parsing), or stratifying evaluation data by semantic complexity (Antoine et al., 2024). Then there is a large area of probing systems with linguistic structures Tenney et al. (2019b); Rama, Beinborn, and Eger (2020); Pimentel et al. (2020b); Starace et al. (2023), also tailored at certain branches of linguistics, e.g., psycholinguistics Gauthier et al. (2020); Ettinger (2020). Moore (2009) has examples of challenges that require linguistic knowledge to unpack, such as structural parallelism and translation of “WH” questions.
In sum, linguistics is crucial for diagnostics of NLP technologies. Besides tracking system progress, it fosters interpretability (section 6), and calls for resource creation (section 3).
Human evaluation and meta-evaluation
The combination of varied human evaluation practices with progress towards highly fluent models has led to diminished reliability of human evaluations: Since generated text has become so fluent, it is harder for humans to distinguish their quality. As a result, not all human evaluation studies are useful measures of model output (Clark et al., 2021; Freitag et al., 2021a), necessitating standardization with regard to design and terminology (Howcroft et al., 2020; Belz, Mille, and Howcroft, 2020; van der Lee et al., 2019). Making use of implicit linguistic knowledge held by native speakers (e.g., judging the fluency of system output) or more explicit linguistic training (e.g. assessing the syntactic diversity of system output), can serve a crucial role in reliably assessing the state of the field (Michael et al., 2016; Callison-Burch et al., 2008).
To design approaches for meta-evaluation (assessing and comparing metrics), expertise in linguistics can prove highly useful, since the challenging linguistic phenomena for models and metrics need to be understood in order to be identified. In a WMT metric shared task, Freitag et al. (2021b) find that most metrics struggle to handle complex semantic phenomena and do not appropriately punish the reversal of sentiment or negation; this is established with test sets that assess specified linguistic features (here: sentiment or negation reversal).
New metrics
Linguistic annotations can help build automatic metrics. Some NLG metrics are built from entailment data [^7] While entailment annotations may come from non-linguists, the notion of textual entailment was inspired by traditions from logic and formal semantics Dagan, Glickman, and Magnini (2006); Bowman et al. (2015). (Chen and Eger, 2023; Xie et al., 2021; Scialom et al., 2021; Steen et al., 2023). Measures for biases, like social biases, have drawn upon the efforts of linguists. For instance, Steen and Markert (2024) propose bias evaluation measures from socio-linguistic lexicons (gender, race) and OntoNotes (Pradhan et al., 2013).
5 Low-Resource Settings
The term “low-resource” has multiple meanings Hedderich et al. (2021). In NLP and ML, we wish to efficiently learn to generalize to new tasks, domains, and languages, including global and historic languages, where feedback and data is sparse or nonexistent.
Processing of global and historic languages
Linguistics is crucial in situations where the amount of language data is limited, e.g., when developing technologies for languages with a limited amount of available recordings or written data. We can improve generalization performance across an array of LLM models by exploiting linguistic features in training (Zhang et al., 2024) or dictionaries for balanced data augmentation (Lu et al., 2024); also with benefits to speech processing (Kim, Jeon, and Lee, 2024). Processing particular types of language/particular languages can also pose challenges that require input from linguistics. For examples, polysynthetic languages such as those spoken in Canada (Gupta and Boulianne, 2020a, b) pose problems for automatic speech recognition (ASR): if ‘word’ is taken as the basic unit for recognition (which makes sense for most languages, but not these ones) the out-of-vocabulary (OOV) rate on new data is extreme.
At first glance, Tanzer et al.’s (2024) study may appear as a potential ‘linguistics-free’ avenue for addressing some extremely low resource-scenarios, since it suggests that LLMs could, to a certain extent, translate a language from one linguist’s grammar book instead of a corpus. But by contrary, the study shows crucial reliance on linguistics: a linguist is required to assess LLM performance, and even the largest LLMs considerably underperform relative to a linguist (unfamiliar with the language) who has read the book. Of course, it is also the case that the fundamental resource was created by a linguist.
Even when more data is available, linguistics can be of value: It can inform data selection in multi-lingual pretraining Ogunremi, Jurafsky, and Manning (2023), or more language-balanced tokenization Creutz and Lagus (2002); Limisiewicz et al. (2024); Fusco et al. (2024). [^8] Some research also suggests that “simply scaling up the number of languages […] in the pretraining is unhelpful” for generalization (Adelani et al., 2024), and “bigger is not always better” Wilcox et al. (2024).